RC Update and Run plan for Aug 31,2020
RC Meeting Information
RC Daily Meetings are at 4:15 PM, remote
- Data taking with PolHe3 longitudinal Kin A/X (SHMS -7.5 GeV 11 degrees, HMS -4.2 GeV 13.50 degrees) until ~7 pm
- Data taking with PolHe3 longitudinal KIN B/Y (SHMS -6.4 GeV 14.5 degrees, HMS -4.2 GeV 16.40 degrees)
- During swing shift - only 24 uA due to RF trips (along with BLM) https://logbooks.jlab.org/entry/3837995
- HallC:p & spin precision & polarization: https://logbooks.jlab.org/entry/3838004
Recent NMR Measurements
- @7:23pm: 42.2% (longitudinal 180 deg) - https://logbooks.jlab.org/entry/3837976
- @5:35am: 42.5% (longitudinal 180 deg) - https://logbooks.jlab.org/entry/3838265
Pass Change & Today's Access
- Simona's Solid Cherenkov work https://logbooks.jlab.org/entry/3838510
- Spectrometer rotation (SHMS --> 8.5 deg & HMS --> 11.7 deg) (Amy/Steve) https://logbooks.jlab.org/entry/3838427
- 3He gas bottle open to prepare reference cell for running https://logbooks.jlab.org/entry/3838431.
- Evacuated and ready to be filled later with hydrogen for 1-pass running. https://logbooks.jlab.org/entry/3838445
- Water purged by compressor (Larry)
- Fast raster reconfiguration for 1-pass running (Mark) https://logbooks.jlab.org/entry/3838507
- Collimator installation (RadCon, JP, techs)
- Target ladder is broken due to interference between the collimator support and target ladder frame.
- The same components and no interference in January.
- Possible shift of the mount support bolts during today's work?
- More panels are taken out to open and check the radiation. Samples are being analyzed by RadCon.
Pending problem(s) / Things to track
- Spare reference cell options/status [Todd, JP]
- Target mounting for Tommy (next cell) [Arun, et al]
- Target mounting for Butterball [Arun, et al]
General Reminders for Running
- Make sure the raster size is correct (4--4.5mm)
- Monitor beam parameters/configuration very often and post an entry like this at the beginning of your shift and when run conditions have changed
- Running on cell targets: raster always ON (4--4.5mm, size will be determined by the scraping checks), orbit locks ON, current ramp rate 1 muA/sec
- Always take 120Hz of EDTM
- Fill the BTA with realistic numbers. If the desired beam current is not delivered prorate your ABUs!
- Change the IHWP at least every 8 hours when running on glass cells. Call Hall A in advance to request. If there are any issues, call the RC
- Note that the autoload of IHWP status in the Start-Run GUI doesn't work anymore so please double check IHWP setting before you accept the GUI value.
Notices in Effect for Shift Crew
- Follow all COVID rules and regulations as well as remind each other. Additional food guild lines:
- All food are removed from the round table. Fruit are now store in the fridge, and individually packaged snacks are store on the shelf below the coffee machine.
- No food near the SL and TO control station. If you would like to eat something, please use the empty round table (used to store food in the past), that is now empty and sanitized.
- New coffee guide line: it is much preferred if you can bring you mug for the coffee. During the shift if one need to drink coffee near the control station, then the coffee mug must have a lid or have some kind of cover.
- Beam positions: 
- Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
- Inform the RC of any access that the Target Team requires. The RC will coordinate with RadCon to make everything run smoothly.
- Please make sure that the orbit locks are enabled and the beam does not drift from its nominal position.
- Ramp rate lock instructions are here and added to checklist
- An new reference Cell refill procedure: New procedure
- Magnet policies
- Spectrometer GUI (
go_magnets) should always be visible on Shift Leader computer, hcdesk3
- Slow ramping procedure of the SHMS, go small steps to the MOL, an great example (1 hour 30 min process)
- HMS Q1 must be ramped down in 50A steps in all cases
- Do not ramp HMS Q2 to zero, only go to 10A and then switch off power.
- New updates to the magnet momentum ramping procedure: https://hallcweb.jlab.org/wiki/index.php/How_to_operate_and_monitor_the_Spectrometer_Magnets
- If any of the Spectrometers magnets trip, take a screenshot of the PSU window and check the hall for smoke or water with the camera before resetting the interlock and post it to hclog
- Spectrometer GUI (
- When rotating the spectrometer, use HCCAM01 to keep an eye on the SOLID Cherenkov test stand when rotating SHMS (in addition to watching the live angle-cam image, etc). The updated rotation instruction can be found: https://hallcweb.jlab.org/wiki/index.php/Spectrometer_Rotation#NOTICES_IN_EFFECT
- New SHMS matrix element
Planned Accelerator Configuration Changes
- Hall C to 1st pass on 8/31 and back to 5th pass on 9/03
- Bluejeans remote connection information:
To join the Meeting: https://bluejeans.com/861439920 To join via Room System: Video Conferencing System: bjn.vc -or-22.214.171.124 Meeting ID : 861439920 To join via phone : 1) Dial: +1.408.740.7256 (US (San Jose)) +1.888.240.2560 (US Toll Free)) +1.408.317.9253 (US (Primary, San Jose)) (see all numbers - http://bluejeans.com/numbers) 2) Enter Conference ID : 861439920